OpenStudy (anonymous):

Which is a solution to the equation 6x + y = –4? A. (1, –10) B. (2, –8) C. (1, 2) D. (2, 8)

OpenStudy (anonymous):

6x + y = –4 plug in choice A. It works ! 6(1) + (-10) = –4 6 - 10) = –4 -4 = –4

OpenStudy (anonymous):

So it's A?

OpenStudy (yanasidlinskiy):

Just like student_basil said. You never know which one works correctly. I will work out B for you and let you see. Just plug it in. 6x+y=-4 6(2) +(-8) = -4 12 + -8=-4 \[4\neq -4\]

OpenStudy (anonymous):

??

OpenStudy (yanasidlinskiy):

And yes! Choice A is correct!

OpenStudy (anonymous):

A carpenter charges a one-time fee of $50 plus $35 for each hour on the job. If n is the number of hours on the job and c is the total charge for the work done, then the variables are related by the equation c = 35n + 50. Which variable is the independent variable? A. n B. c C. There is no relationship between the variables.

OpenStudy (anonymous):

A carpenter charges a one-time fee of $50 plus $35 for each hour on the job. If n is the number of hours on the job and c is the total charge for the work done, then the variables are related by the equation c = 35n + 50. Which variable is the independent variable? A. n B. c C. There is no relationship between the variables.

OpenStudy (anonymous):

@thomaster